OpenStudy (chillhill):

Find a linear function f(x) = ax+b for which f(2)=1 and f(-3) =11

OpenStudy (jango_in_dtown):

so we have to compute f(x)=y=ax+b see by the given data (2,1) and (-3,11) are the points on the line y=ax+b

OpenStudy (jango_in_dtown):

so compute the slope first which is (y2-y1)/(x2-x1)=(11-2)/(-3-2)=10/(-5)=-2 so a=slope=-2

OpenStudy (jango_in_dtown):

now the equation takes the form y=-2x+b put x=2 and y=1 (since (2,1) satisfies the line) and you have 1=-2.2+b implies b=1+4=5

OpenStudy (jango_in_dtown):

so the final equation is y=f(x)=-2x+5

OpenStudy (jango_in_dtown):

@chillhill

OpenStudy (chillhill):

Thank you so much for your help